Nick has $2.95 in quarters and dimes. If he has 19 coins in all, how many of each does he have?

Answer by Cromlix(4381)

Quarters = x
Dimes = y
x + y = 19.........1
25x 10y = 295......2
Multiply (1) by 10
10x + 10y = 190.....3
25x + 10y = 295.....2
Subtract (3) from (2)
15x = 105
x = 7
Substitute x = 7 in
Eq (1)
x + y = 19
7 + y = 19
y = 19 - 7
y = 12
Nick had 7 quarters and 12 dimes.
Hope this helps.
:-)